OpenStudy (graceanne155):

You spike a volley ball on the ground so it rebounds with a velocity = 0 m/s i, 5 m/s j. How long will it be in the air for?

OpenStudy (irishboy123):

that means that \[\vec v_o = \left(\begin{matrix}0 \\ 5\end{matrix}\right)\] so use your equation of motion i'd suggest \(v = u - g t \) use symmetry: it's t for up and then same t for back down